# Large Language Models for Cyber Security: A Systematic Literature Review

### v4

@@ Abstract.
Through our analysis, we identify several key findings. First, we observe that LLMs are being applied to an expanding range of cybersecurity tasks, including vulnerability detection, malware analysis, network intrusion detection, and phishing detection. Second, we analyze how different LLM architectures (such as encoder-only, encoder-decoder, and decoder-only) are utilized across various security domains and over time, identifying key application trends in LLM4Security. Third, we identify several increasingly sophisticated techniques for adapting LLMs to specific cybersecurity domains, such as advanced fine-tuning, prompt engineering, and external augmentation strategies. Fourth, we find that the datasets used for training and evaluating LLMs in these tasks are often limited in size and diversity, highlighting the need for more comprehensive datasets and the emerging use of LLMs for data augmentation. Finally, we discuss the main challenges and opportunities for future research in LLM4Security, including the need for more interpretable and robust models, the importance of addressing data security and the inherent security risks of LLMs themselves, and the potential for leveraging LLMs for proactive defense and threat hunting.
@@ 1. Introduction
In particular, one important domain where LLMs are beginning to show promise is cybersecurity. With the growing volume and sophistication of cyber threats, there is an urgent need for intelligent systems that can automatically detect vulnerabilities, analyze malware, and respond to attacks (Chen et al., 2023c; Bilot et al., 2023; Montes et al., 2021). Recent research has explored the application of LLMs across a wide range of cybersecurity tasks, i.e., LLM4Security hereafter. In the domain of software security, LLMs have been used for detecting vulnerabilities from natural language descriptions and source code, as well as generating security-related code, such as patches and exploits. These models have shown high accuracy in identifying vulnerable code snippets and generating effective patches for common types of vulnerabilities (Chow et al., 2023; Fu et al., 2022; Charalambous et al., 2023). Beyond code-level analysis, LLMs have also been applied to understand and analyze higher-level security artifacts, such as security policies and privacy policies, helping to classify documents and detect potential violations (Hartvigsen et al., 2022; Mets et al., 2023). In the realm of network security, LLMs have demonstrated the ability to detect and classify various types of attacks from network traffic data, including DDoS attacks, port scanning, and botnet traffic (Alkhatib et al., 2022; Moskal et al., 2023; Ali and Kostakos, 2023). Malware analysis is another key area where LLMs are showing promise, with models being used to classify malware families based on textual analysis reports and behavioral descriptions, as well as detecting malicious domains and URLs (Joyce et al., 2023; Liu et al., 2023e). LLMs have also been employed in the field of social engineering to detect and defend against phishing attacks by analyzing email contents and identifying deceptive language patterns (Jamal et al., 2024; Ranade et al., 2021). Moreover, researchers are exploring the use of LLMs to enhance the robustness and resilience of security systems themselves, by generating adversarial examples for testing the robustness of security classifiers and simulating realistic attack scenarios for training and evaluation purposes (Temara, 2023; Charan et al., 2023; Saha et al., 2024). These diverse applications demonstrate the significant potential of LLMs to improve the efficiency and effectiveness of cybersecurity practices by processing and extracting insights from large amounts of unstructured text, learning patterns from vast datasets, and generating relevant examples for testing and training purposes.
@@
To address these limitations and provide an in-depth understanding of the state-of-the-art in LLM4Security, we conduct a systematic and extensive survey of the literature. By comprehensively collecting 47,135 relevant papers and systematically analyzing 185 papers from top security and software engineering venues, our survey aims to provide a holistic view of how LLMs are being applied to solve diverse problems across the cybersecurity domain. In addition to identifying the types of tasks that LLMs are being used for, we also examine the specific datasets, preprocessing techniques, and domain adaptation methods employed in each case. This enables us to provide a more nuanced analysis of the strengths and limitations of different approaches, and to identify the most promising directions for future research. Specifically, we focus on answering four key research questions (RQs):
@@
For each research question, we provide a fine-grained analysis of the approaches, datasets, and evaluation methodologies used in the surveyed papers. We identify common themes and categorize the papers along different dimensions to provide a structured overview of the landscape. Furthermore, we highlight the key challenges and limitations of current approaches to guide future research towards addressing the gaps. We believe our survey can serve as a valuable resource for researchers working at the intersection of NLP, AI, and cybersecurity. The contributions of this work are summarized as follows:
  • We conduct a comprehensive Systematic Literature Review (SLR) to investigate the latest research on LLM4Security, providing a mapping of the current landscape. Our search covers an extensive number of over 47,135 papers. With further quality-based and relevance-based filtering, we retain 185 papers for later detailed review.
  • We formulate four key RQs to understand various aspects of LLM application in security in each distinct dimension, including the types of LLMs used, security tasks facilitated, domain specification techniques, and differences in data collection and pre-processing.
  • We analyze the distribution of the 185 selected papers across venues and over time, and categorizes the characteristics of mainstream LLMs employed in the security domain.
@@ 2.3.1. Coarse-Grained Inclusion and Exclusion Criteria
To select relevant papers for our research questions, we defined four inclusion criteria and eight exclusion criteria (as listed in Table 2) for the coarse-grained paper selection process. Among them, In#1, Ex#1, Ex#2, and Ex#3 were automatically filtered based on the keywords, duplication status, length, and publication venue of the papers. The remaining inclusion criteria (In#2~4) and exclusion criteria (Ex#4~8) were manually applied by inspecting the topic and content of each paper. Specifically, the criteria of In#1 retained 10,302 papers whose titles and abstracts contained a pair of the identified search keywords. Subsequently, Ex#1 filtered out 797 duplicate or multi-version papers from the same authors with little difference. Next, the automated filtering criteria Ex#2 was applied to exclude short papers, tool demos, keynotes, editorials, books, theses, workshop papers, or poster papers, resulting in 6,507 papers being removed. The remaining papers were then screened based on the criteria Ex#3, which retained 622 full research papers published in the identified venues or as preprints on arXiv. The remaining inclusion and exclusion criteria (In#2~4, Ex#4~8) were then manually applied to the titles and abstracts of these 622 papers, in order to determine their relevance to the research topic. Three researchers independently applied the inclusion and exclusion criteria to the titles and abstracts. Disagreements were resolved through discussion and consensus. After this manual inspection stage, 236 papers were included for further fine-grained full-text quality assessment.
@@ 3. RQ1: What types of security tasks have been facilitated by LLM-based approaches?
This section delves into the detailed examination of LLM utilization across diverse security domains. We have classified them into six primary domains, aligning with the themes of the collected papers: software and system security, network security, information and content security, hardware security, and blockchain security, totaling 185 papers. Figure 4 visually depicts the distribution of LLMs within these six domains. Additionally, Table 4 offers a comprehensive breakdown of research detailing specific security tasks addressed through LLM application.
@@ 3.2. Application of LLMs in Software and System Security
Bug repair. LLMs possess robust code generation capabilities, and their utilization in engineering for code generation can significantly enhance efficiency. However, code produced by LLMs often carries increased security risks, such as bugs and vulnerabilities (Perry et al., 2023). These program bugs can lead to persistent security vulnerabilities. Hence, automating the process of bug fixing is imperative, involving the use of automation technology to analyze flawed code and generate accurate patches to rectify identified issues. LLMs like CodeBERT (Le-Cong et al., 2023; Xia and Zhang, 2022; Huang et al., 2023; Zhang et al., 2023b), CodeT5 (Wang et al., 2023b; Tang et al., 2023; Huang et al., 2023; Liu et al., 2024c), Codex (Jin et al., 2023; Fan et al., 2023; Xia and Zhang, 2023; Chow et al., 2024; Tang et al., 2024), LLaMa (Tang et al., 2023), CodeLLaMa (Olausson et al., 2024; Silva et al., 2024) , CodeGEN (Xia and Zhang, 2023), UniXcoder (Zhang et al., 2023b), T5 (Yuan et al., 2022), PLBART (Huang et al., 2023), Starcoder (Zhao et al., 2024) and GPT Series (Zhang et al., 2023c; Olausson et al., 2024; Zhang et al., 2023b; Xia and Zhang, 2024, 2023; Tang et al., 2023; Zhang et al., 2023f; Yang et al., 2024b; Yin et al., 2024b; Zhang et al., 2024c; Huang et al., 2024; Feng and Chen, 2024; Zhong et al., 2024; Wang et al., 2024c) have showcased effectiveness in generating syntactically accurate and contextually relevant code. This includes frameworks with encoder-decoder architecture like Repilot (Wei et al., 2023c), tailored specifically for producing repair patches. Utilizing LLMs for program repair can achieve competitive performance in producing patches for various types of errors and defects (Xia and Zhang, 2024). These models effectively capture the underlying semantics and dependencies in code, resulting in precise and efficient patches. Moreover, fine-tuning LLMs on specific code repair datasets can further improve their ability to generate high-quality patches for real-world software projects. Integrating LLMs into program repair not only speeds up the error-fixing process but also allows software developers to focus on more complex tasks, thereby enhancing the reliability and maintainability of the software (Xia and Zhang, 2023). As demonstrated in the case of ChatGPT, notably enhances the accuracy of program repairs when integrated with interactive feedback loops (Xia and Zhang, 2023). This iterative process of patch generation and validation fosters a nuanced comprehension of software semantics, thereby resulting in more impactful fixes. Complementary to generating fixes, evaluating the capabilities of LLM-driven code agents requires realistic benchmarks; SWT-Bench, for example, provides a framework based on real-world data to jointly assess test generation and fix validation performance (Mündler et al., 2024). By integrating domain-specific knowledge and technologies with the capabilities of LLMs, their performance is further enhanced. Custom prompts, fine-tuning for specific tasks, retrieving external data, and utilizing static analysis tools (Xia et al., 2023; Tang et al., 2023; Zhang et al., 2023a; Fu et al., 2022; Jin et al., 2023) significantly improve the effectiveness of bug fixes driven by LLMs.
@@
Reverse engineering and binary analysis. Reverse engineering is the process of attempting to understand how existing artifacts work, whether for malicious purposes or defensive purposes, and it holds significant security implications. The capability of LLMs to recognize software functionality and extract important information enables them to perform certain reverse engineering steps (Pearce et al., 2022). For example, Xu et al. (Xu et al., 2023) achieved recovery of variable names from binary files by propagating LLMs query results through multiple rounds. Techniques like GENNM further advance this by using generative LLMs with context-aware fine-tuning, preference optimization, and iterative inference to recover names from stripped binaries (Xu et al., 2025). Armengol-Estapé et al. (Armengol-Estapé et al., 2024) combined type inference engines with LLMs to perform disassembly of executable files and generate program source code. LLMs can also be used to assist in binary program analysis. Sun et al. (Sun et al., 2023) proposed DexBert for characterizing Android system binary bytecode. Pei et al. (Pei et al., 2024) preserved the semantic symmetry of code based on group theory, resulting in their binary analysis framework SYMC demonstrating outstanding generalization and robustness in various binary analysis tasks. Song et al. (Song et al., 2022) utilized LLMs to address authorship analysis issues in software engineering, effectively applying them to real-world APT malicious software for organization-level verification. Some studies (Hu et al., 2024a; She et al., 2024; Fang et al., 2024b) apply LLMs to enhance the readability and usability of decompiler outputs, including decompiling WebAssembly into higher-level languages like C/C++, thereby assisting reverse engineers in better understanding binary files.

### v5

@@ Abstract.
Through our analysis, we identify several key findings. First, we observe that LLMs are being applied to an expanding range of cybersecurity tasks, including vulnerability detection, malware analysis, and network intrusion detection. Second, we analyze application trends of different LLM architectures (such as encoder-only, encoder-decoder, and decoder-only) across security domains. Third, we identify increasingly sophisticated techniques for adapting LLMs to cybersecurity, such as advanced fine-tuning, prompt engineering, and external augmentation strategies. A significant emerging trend is the use of LLM-based autonomous agents, which represent a paradigm shift from single-task execution to orchestrating complex, multi-step security workflows. Furthermore, we find that the datasets used for training and evaluating LLMs are often limited, highlighting the need for more comprehensive datasets and the use of LLMs for data augmentation. Finally, we discuss the main challenges and opportunities for future research, including the need for more interpretable models, addressing the inherent security risks of LLMs, and their potential for proactive defense.
@@ 1. Introduction
In particular, one important domain where LLMs are beginning to show promise is cybersecurity. With the growing volume and sophistication of cyber threats, there is an urgent need for intelligent systems that can automatically detect vulnerabilities, analyze malware, and respond to attacks (Chen et al., 2023c; Bilot et al., 2023; Montes et al., 2021). Recent research has explored the application of LLMs across a wide range of cybersecurity tasks, i.e., LLM4Security hereafter. In the domain of software security, LLMs have been used for detecting vulnerabilities from natural language descriptions and source code, as well as generating security-related code, such as patches and exploits. These models have shown high accuracy in identifying vulnerable code snippets and generating effective patches for common types of vulnerabilities (Chow et al., 2023; Fu et al., 2022; Charalambous et al., 2023). Beyond code-level analysis, LLMs have also been applied to understand and analyze higher-level security artifacts, such as security policies and privacy policies, helping to classify documents and detect potential violations (Hartvigsen et al., 2022; Mets et al., 2023). In the realm of network security, LLMs have demonstrated the ability to detect and classify various types of attacks from network traffic data, including DDoS attacks, port scanning, and botnet traffic (Alkhatib et al., 2022; Moskal et al., 2023; Ali and Kostakos, 2023). Malware analysis is another key area where LLMs are showing promise, with models being used to classify malware families based on textual analysis reports and behavioral descriptions, as well as detecting malicious domains and URLs (Joyce et al., 2023; Liu et al., 2023e).
LLMs have also been employed in the field of social engineering to detect and defend against phishing attacks by analyzing email contents and identifying deceptive language patterns (Jamal et al., 2024; Ranade et al., 2021). Moreover, researchers are exploring the use of LLMs to enhance the robustness and resilience of security systems themselves, by generating adversarial examples for testing the robustness of security classifiers and simulating realistic attack scenarios for training and evaluation purposes (Temara, 2023; Charan et al., 2023; Saha et al., 2024). These diverse applications demonstrate the significant potential of LLMs to improve the efficiency and effectiveness of cybersecurity practices by processing and extracting insights from large amounts of unstructured text, learning patterns from vast datasets, and generating relevant examples for testing and training purposes.
@@
Scope. To address these limitations and provide an in-depth understanding of the state-of-the-art in LLM4Security, we conduct a systematic and extensive survey of the literature. By comprehensively collecting 47,135 relevant papers and systematically analyzing 185 papers from top security and software engineering venues, our survey aims to provide a holistic view of how LLMs are being applied to solve diverse problems across the cybersecurity domain. In addition to identifying the types of tasks that LLMs are being used for, we also examine the specific datasets, preprocessing techniques, and domain adaptation methods employed in each case. This enables us to provide a more nuanced analysis of the strengths and limitations of different approaches, and to identify the most promising directions for future research. Specifically, we focus on answering four key research questions (RQs):
@@
For each research question, we provide a fine-grained analysis of the approaches, datasets, and evaluation methodologies used in the surveyed papers. We identify common themes and categorize the papers along different dimensions to provide a structured overview of the landscape. Furthermore, we highlight the key challenges and limitations of current approaches to guide future research towards addressing the gaps.
Significance. A key significance of this survey lies in its deep-seated connection to Software Engineering (SE) community. This connection is not incidental; our systematic selection process deliberately includes premier SE venues (e.g., ICSE, FSE, ASE, etc.), and our analysis reveals that the vast majority of research (63%) falls under the umbrella of software and system security. The surveyed applications of LLMs are fundamentally addressing core challenges within the secure software development lifecycle, ranging from vulnerability detection and automated program repair to fuzzing and secure code generation. More profoundly, we observe that the methodologies employed in LLM4Security represent an evolution of established SE research paradigms (Hou et al., 2024). For instance, the generative capabilities of LLMs are pushing the boundaries of classic SE tasks like automated testing and program repair, while the reliance on curated code corpora and the novel use of LLMs for data augmentation align directly with the principles of data-driven software engineering. Therefore, this survey not only maps a subfield of cybersecurity but also chronicles a significant, ongoing transformation within software engineering itself.
Contributions. The contributions of this work are summarized as follows:
  • We provide the first comprehensive and structured map of the LLM4Security landscape, synthesized from a Systematic Literature Review (SLR) of 185 high-quality papers selected from an initial pool of over 47,000 candidates.
  • We deliver a multi-dimensional analysis of the field by systematically investigating four crucial aspects across the selected studies: the security tasks being automated, the specific LLM architectures being employed, the domain-specific techniques (e.g., fine-tuning and external augmentation) used for adaptation, and the data-centric practices for training and evaluation.
  • We synthesize and present key insights into how LLMs are reshaping the Secure-Software-Engineering lifecycle. Our findings reveal a paradigm shift from mere defect detection towards automated generation and repair; the use of LLMs for data augmentation to tackle data scarcity; and the emergence of hybrid workflows where LLMs orchestrate traditional SE tools, heralding a new era of security automation.
@@ 2.3.1. Coarse-Grained Inclusion and Exclusion Criteria
To select relevant papers for our research questions, we defined four inclusion criteria and eight exclusion criteria (as listed in Table 2) for the coarse-grained paper selection process. Among them, In#1, Ex#1, Ex#2, and Ex#3 were automatically filtered based on the keywords, duplication status, length, and publication venue of the papers. The remaining inclusion criteria (In#2˜4) and exclusion criteria (Ex#4˜8) were manually applied by inspecting the topic and content of each paper. Specifically, the criteria of In#1 retained 10,302 papers whose titles and abstracts contained a pair of the identified search keywords. Subsequently, Ex#1 filtered out 797 duplicate or multi-version papers from the same authors with little difference. Next, the automated filtering criteria Ex#2 was applied to exclude short papers, tool demos, keynotes, editorials, books, theses, workshop papers, or poster papers, resulting in 6,507 papers being removed. The remaining papers were then screened based on the criteria Ex#3, which retained 622 full research papers published in the identified venues or as preprints on arXiv. The remaining inclusion and exclusion criteria (In#2˜4, Ex#4˜8) were then manually applied to the titles and abstracts of these 622 papers, in order to determine their relevance to the research topic. Three researchers independently applied the inclusion and exclusion criteria to the titles and abstracts. Disagreements were resolved through discussion and consensus. After this manual inspection stage, 236 papers were included for further fine-grained full-text quality assessment.
@@ 3. RQ1: What types of security tasks have been facilitated by LLM-based approaches?
This section delves into the detailed examination of LLM utilization across diverse security domains. We have classified them into six primary domains, aligning with the themes of the collected papers: software and system security, network security, information and content security, hardware security, and blockchain security, totaling 185 papers. Figure 4 visually depicts the distribution of LLMs within these six domains. Additionally, Table 4 offers a comprehensive breakdown of research detailing specific security tasks addressed through LLM application. We also analyze the recent and significant trend of employing LLM-based autonomous agents, which represent a paradigm shift from performing isolated tasks to orchestrating complex security workflows.
@@ 3.2. Application of LLMs in Software and System Security
Bug repair. LLMs possess robust code generation capabilities, and their utilization in engineering for code generation can significantly enhance efficiency. However, code produced by LLMs often carries increased security risks, such as bugs and vulnerabilities (Perry et al., 2023). These program bugs can lead to persistent security vulnerabilities. Hence, automating the process of bug fixing is imperative, involving the use of automation technology to analyze flawed code and generate accurate patches to rectify identified issues. LLMs like CodeBERT (Le-Cong et al., 2023; Xia and Zhang, 2022; Huang et al., 2023; Zhang et al., 2023b), CodeT5 (Wang et al., 2023b; Tang et al., 2023; Huang et al., 2023; Liu et al., 2024c), Codex (Jin et al., 2023; Fan et al., 2023; Xia and Zhang, 2023; Chow et al., 2024; Tang et al., 2024), LLaMa (Tang et al., 2023), CodeLLaMa (Olausson et al., 2024; Silva et al., 2024) , CodeGEN (Xia and Zhang, 2023), UniXcoder (Zhang et al., 2023b), T5 (Yuan et al., 2022), PLBART (Huang et al., 2023), Starcoder (Zhao et al., 2024) and GPT Series (Zhang et al., 2023c; Olausson et al., 2024; Zhang et al., 2023b; Xia and Zhang, 2024, 2023; Tang et al., 2023; Zhang et al., 2023f; Yang et al., 2024b; Yin et al., 2024b; Zhang et al., 2024c; Huang et al., 2024b; Feng and Chen, 2024; Zhong et al., 2024; Wang et al., 2024c) have showcased effectiveness in generating syntactically accurate and contextually relevant code. This includes frameworks with encoder-decoder architecture like Repilot (Wei et al., 2023c), tailored specifically for producing repair patches. Utilizing LLMs for program repair can achieve competitive performance in producing patches for various types of errors and defects (Xia and Zhang, 2024). These models effectively capture the underlying semantics and dependencies in code, resulting in precise and efficient patches. Moreover, fine-tuning LLMs on specific code repair datasets can further improve their ability to generate high-quality patches for real-world software projects. Integrating LLMs into program repair not only speeds up the error-fixing process but also allows software developers to focus on more complex tasks, thereby enhancing the reliability and maintainability of the software (Xia and Zhang, 2023). As demonstrated in the case of ChatGPT, notably enhances the accuracy of program repairs when integrated with interactive feedback loops (Xia and Zhang, 2023). This iterative process of patch generation and validation fosters a nuanced comprehension of software semantics, thereby resulting in more impactful fixes. Complementary to generating fixes, evaluating the capabilities of LLM-driven code agents requires realistic benchmarks; SWT-Bench, for example, provides a framework based on real-world data to jointly assess test generation and fix validation performance (Mündler et al., 2024). By integrating domain-specific knowledge and technologies with the capabilities of LLMs, their performance is further enhanced. Custom prompts, fine-tuning for specific tasks, retrieving external data, and utilizing static analysis tools (Xia et al., 2023; Tang et al., 2023; Zhang et al., 2023a; Fu et al., 2022; Jin et al., 2023) significantly improve the effectiveness of bug fixes driven by LLMs.
@@
Reverse engineering and binary analysis. Reverse engineering is the process of attempting to understand how existing artifacts work, whether for malicious purposes or defensive purposes, and it holds significant security implications. The capability of LLMs to recognize software functionality and extract important information enables them to perform certain reverse engineering steps (Pearce et al., 2022). For example, Xu et al. (Xu et al., 2023) achieved recovery of variable names from binary files by propagating LLMs query results through multiple rounds. Techniques like GENNM further advance this by using generative LLMs with context-aware fine-tuning, preference optimization, and iterative inference to recover names from stripped binaries (Xu et al., 2025b). Armengol-Estapé et al. (Armengol-Estapé et al., 2024) combined type inference engines with LLMs to perform disassembly of executable files and generate program source code. LLMs can also be used to assist in binary program analysis. Sun et al. (Sun et al., 2023) proposed DexBert for characterizing Android system binary bytecode. Pei et al. (Pei et al., 2024) preserved the semantic symmetry of code based on group theory, resulting in their binary analysis framework SYMC demonstrating outstanding generalization and robustness in various binary analysis tasks. Song et al. (Song et al., 2022) utilized LLMs to address authorship analysis issues in software engineering, effectively applying them to real-world APT malicious software for organization-level verification. Some studies (Hu et al., 2024a; She et al., 2024; Fang et al., 2024b) apply LLMs to enhance the readability and usability of decompiler outputs, including decompiling WebAssembly into higher-level languages like C/C++, thereby assisting reverse engineers in better understanding binary files.
@@ 3.5. Application of LLMs in Blockchain Security
3.6. Application of LLM Agents in Security
Beyond single-task applications, a significant recent trend is the development of LLM-based autonomous agents for security. These agents are sophisticated systems built around an LLM core, endowed with capabilities for autonomous reasoning, planning, memory, and the use of external tools to accomplish complex, multi-step goals (Xu et al., 2025a; Luo et al., 2025). This paradigm shifts from using LLMs as simple query-response tools to employing them as autonomous problem-solvers that can interact with environments and other agents to tackle dynamic security challenges. A predominant architecture emerging is the Multi-Agent System (MAS), where specialized agents collaborate, debate, or operate in a hierarchy to achieve a common objective.
Network Security. In network security, agent-based frameworks are revolutionizing complex offensive tasks. For penetration testing, multi-agent systems like VulnBot (Kong et al., 2025) and hierarchical teams of agents (Zhu et al., 2024) automate the entire attack lifecycle. These systems typically feature a planning agent that decomposes the goal, specialized agents that execute tasks like reconnaissance or tool use (e.g., using ‘nmap‘ or ‘Metasploit‘), and a reporting agent that synthesizes the findings. This collaborative approach has proven effective in exploiting even zero-day vulnerabilities. For protocol analysis, agents like RFCScan (Zheng et al., 2025) employ a dual-agent structure, with an ”indexing agent” processing RFC documents and a ”detection agent” querying this indexed knowledge to find functional bugs.
Software and System Security. This domain has seen a surge in agentic systems that address the full lifecycle of a vulnerability. For vulnerability detection and exploitation, frameworks like FaultLine (Nitin et al., 2025) establish an agentic workflow to automate the generation of proof-of-vulnerability exploits, moving beyond simple detection to active confirmation. For vulnerability repair, multi-agent systems like AutoPatch (Seo et al., 2025) create a pipeline of specialized agents, including a ”Vulnerability Verifier” and a ”Code Patcher,” that work together to automatically fix insecure code generated by LLMs. In proactive defense, SCGAgent (Saul et al., 2025) functions as an autonomous agent that leverages security guidelines and generates unit tests to ensure newly written code is both functional and secure. For malware analysis, agents like LAMD (Qian et al., 2025) emulate a human analyst’s workflow by first using backward program slicing—a classic program analysis technique—to isolate security-critical code regions. This focused context is then fed into a tier-wise reasoning process, allowing the agent to analyze behavior from individual instructions to overall intent and autonomously detect malicious patterns in Android applications. Furthermore, agents are being developed for the complex task of reverse engineering and binary analysis. ReCopilot (Chen et al., 2025) is an expert agent trained specifically for binary analysis tasks like function name recovery from decompiled code, outperforming general-purpose models. For end-to-end analysis, Vul-BinLLM (Hussain et al., 2025) detects vulnerabilities directly in compiled binaries by using an optimized decompiler and a function analysis queue to manage the LLM’s context window, allowing it to analyze large, real-world binaries while maintaining overall context. A key insight from this research is that the most successful software security agents are not replacing classical tools but orchestrating them; the LLM acts as the intelligent ”brain” while requiring the precise ”eyes and ears” of static analyzers, debuggers, and program slicers to operate effectively.
Information and Content Security. To combat deception, researchers have developed innovative debate-driven multi-agent systems. In phishing detection, frameworks like PhishDebate (Li et al., 2025a) and similar debate-oriented models (Li et al., 2025b) utilize a team of agents. Each agent is assigned a specific role, such as analyzing URL structures, scrutinizing textual content for social engineering tactics, or performing screenshot analysis. These agents then debate their findings to reach a collective, more robust conclusion, significantly reducing false positives and outperforming single-agent approaches. Beyond deception detection, the capabilities of LLM agents also introduce novel challenges and applications. For access control, as agents gain autonomy, the risk of privilege escalation becomes a paramount concern. To address this, systems like Prompt Flow Integrity (Kim et al., 2025) have been proposed to enforce the principle of least privilege by identifying untrusted data sources and validating unsafe data flows.
Hardware Security. Hardware security verification has traditionally been a manual, labor-intensive process, making it a prime candidate for automation through LLM agents that can learn the complex semantics of hardware description languages. Frameworks like BugWhisperer (Tarek et al., 2025) are leading examples, applying agent-like automation to identify vulnerabilities in SoC designs. It utilizes a specialized LLM, fine-tuned on a comprehensive hardware vulnerability database, to automatically detect security flaws at the Register-Transfer Level (RTL). Other research directions include RTL-Breaker (Mankali et al., 2025), which assesses LLM security against backdoor attacks during HDL generation.
Blockchain Security. The immutable and high-value nature of blockchain systems makes smart contract auditing a critical domain where agents show immense potential. The A1 agentic system (Gervais and Zhou, 2025) exemplifies an execution-grounded approach, transforming a general LLM into an end-to-end exploit generator. The agent hypothesizes a vulnerability, tests an exploit against a forked blockchain state, and uses the deterministic execution feedback to refine its attack, allowing it to overcome hallucinations and discover complex exploits. Other advanced frameworks include Smartify (Karanjai et al., 2025), a multi-agent system for detecting and repairing vulnerabilities, and SmartAuditFlow (Wei et al., 2025), which uses RAG and adaptive planning to improve audit reliability. These systems show that for formal domains like blockchain, the most promising path for agents involves building tight feedback loops between the LLM’s hypothesis generation and a deterministic execution harness.